OpenStudy (anonymous):

According to the reaction shown, how many moles of oxygen gas are necessary to produce 24.0 g of Li2O? 4Li(s)+O2(s)->2Li2O(s) Any help would be appreciated. If someone could walk me through the entire thing as if I know nothing about chemistry, that would help because I'm completely lost in this class right now.

OpenStudy (anonymous):

Alright. First step, we need the amount of \(Li_2O\). To calculate the amount of \(Li_2O\), we use the formula \(\large n_{Li_2O}=\frac{mass}{molar~mass}\). Next the equation is \(4Li~~~~~~~+~~~~~~~O_2~~~~~~~\rightarrow~~~~~~~2Li_2O\) 1 moles 2 moles ? moles \(n\) moles Therefore, \(\Large n_{O_2}=\frac{n_{Li_2O} \times 1}{2}=\frac{n_{Li_2O}}{2}\). Plug \(n\) calculated above in. You will get \(n_{O_2}\)

OpenStudy (vera_ewing):

@JFraser Is the answer to this one 1.6 mol?

OpenStudy (jfraser):

it would be 1.6 moles of lithium, but the question asks about oxygen gas. your last conversion is upside-down